Calculate the concentration of the vinegar used if 75cm3 vinegar reacts completely with 60cm3 of NaHCO3 solution of concentration 0.1mol.dm3 .the equation for the reaction
CH3COOH+NaHCO3-CH3COONa+H20+CO2
1
Expert's answer
2019-05-10T08:05:19-0400

As 1 dm3 = 1 l, the molarity of NaHCO3 is c = 0.1 mol/l.

Let`s convert cm3 into l.

75 cm3 = 0.075 l; 60 cm3 = 0.06 l;

Now, we will count the amount of NaHCO3.

n(NaHCO3) = c * V = 0.1 mol/l * 0.06 l = 0.006 mol.

n(CH3COOH) = n(NaHCO3) = 0.006 mol.

c(CH3COOH) = n/V = 0.006 mol / 0.075 l = 0.08 mol/l = 0.08 mol/dm3.